Circular No. 6527

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996cb IN NGC 3510 Q. Qiao, W. Li, Y. Qiu, and J. Hu, Beijing Astronomical Observatory

(BAO), report their independent discovery of SN 1996cb; their unfiltered CCD images yield the following magnitudes: Nov. 29 UT, [19.0: (limiting magnitude of CCD frame); Dec. 18.92, ∼ 15.9; 19.77, 15.6. A lowresolution, high-quality spectrogram (0.48 nm/pixel, range 400–800 nm) taken with the 2.16-m BAO telescope by D. W. Xu, J. M. Bai, G. W. Cha, and Q. Y. Qiao on Dec. 19.85 confirms SN 1996cb to be a type-II supernova (cf. IAUC 6524); the Balmer lines show strong P-Cyg profiles with expansion velocities of 12 600 (Hα), 7850 (Hβ), and 7210 km/s (Hδ); Na I 589.3-nm and He I 587.6-nm lines also show P-Cyg profiles with expansion velocities of 6670 km/s. The spectrum is similar to that of SN 1996W, two weeks past maximum (on Apr. 29).

NOVA AQUILAE 1995 A. Retter, E. M. Leibowitz, and O. Kovo-Kariti, Wise Observatory,

Tel Aviv University, write: “We performed CCD photometry of N Aql 1995 through an I filter during 3 nights in 1995 and 18 nights in 1996. Time-series analysis of the 1996 data reveals two periodicities in the light curve: 0.060054 ± 0.000004 (which appears with its first harmonic) and 0.2558 ± 0.0001 day, both with full amplitudes of ∼ 0.015 ± 0.005 mag. We cannot rule out that the 1-day aliases at 0.0639 and 0.2037 or 0.3438 day are the true periods. We suggest an intermediate-polar model as an interpretation for the two periods, with the longer one as the orbital period of the binary system.”

COMET C/1996 P2 (RUSSELL-WATSON) Ephemeris extension to IAUC 6488 (orbital elements MPC 28557):

1996 /97 α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Dec. 23 1^{h}00.^{m}63 −26^{o}22.^{′}8 3.744 3.886 90.^{o}9 14.^{o}7 14.8
Jan. 2 1 05.18 −24 33.7 3.957 3.971 83.7 14.2 15.0
12 1 10.80 −22 45.8 4.170 4.057 76.6 13.6 15.2
22 1 17.31 −21 00.3 4.382 4.143 69.6 12.9 15.4
Feb. 1 1 24.52 −19 18.3 4.588 4.229 62.8 12.0 15.6
11 1 32.32 −17 40.6 4.788 4.315 56.1 10.9 15.8
21 1 40.57 −16 07.8 4.977 4.400 49.5  9.8 15.9
Mar. 3 1 49.16 −14 40.3 5.156 4.486 43.3  8.7 16.1
1996 December 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green